OpenStudy (anonymous):

Find the value of x for a to be parallel to b. m<1 = 117 and m<2 = 6x – 12. 2/25??

OpenStudy (anonymous):

What is this m<1, does that mean slope of the first?

OpenStudy (anonymous):

HERES THE SNAPSHOT

OpenStudy (accessdenied):

m<1 and m<2 would be supplementary. 117 + (6x - 12) = 180

OpenStudy (anonymous):

Right......but did you checkout my attachment?

OpenStudy (accessdenied):

Yes. I made that statement based on what I saw in your attachment.

OpenStudy (anonymous):

may you help me out with it?

OpenStudy (accessdenied):

That was my help. You'd solve for x in that equation 117 + (6x - 12) = 180 117 + (6x - 12) = 180 (117 - 12) + 6x = 180 105 + 6x = 180 6x = 75 x = 75 / 6 greatest common factor of 75 and 6 is 3 x = 25(3) / 2(3) = 25/2

OpenStudy (anonymous):

ohhh lol i had it the other away around..thanks!